The distance from the nadir to the base of the building is 3.42 in. and to the top of the same building is 3.60 in. The flying height above sea level was 11,400 feet and the ground elevation of the base of the building is 1400 feet. Calculate the height of the building.
1
Expert's answer
2019-10-16T09:31:00-0400

"H = \frac{(3.6-3.4)(11400-1400) }{3.6}=555.56 feet"


"H = \frac{(n_{t} - n_{base})(S_{fh} - S_{base}) }{n_{t}}"

Where nt -is distance from the nadir to the top of the building, nbase - is distance from the nadir to the base of the building, Sfh - is flying height above sea level, Sbase- the ground elevation of the base of the building